a particle moves with a velocity v(t)=(1/2)kt^2 along a straight line,where is a positive constant. Find the average speed of particle in time t 
 

Grade:11

2 Answers

Vikas TU
4 years ago
V = ½ Kt^2
S is given by integral V . dt 
Put value of v nd solve 
S = Kt^3/6
Average velocity s given by 
Distance travelled / T
=> Kt^2/6 
Arun
4 years ago
Velocity = kt^2 /2
 
distance travelled = Integration of v w.r.t
 = k t^3 /6
 
hence average speed = k t^3 /6 / t = k t^2 /6